# Generalized Nash Equilibrium Solutions in Dynamic Games With Shared Constraints

### v2

@@ Abstract
In dynamic games with shared constraints, Generalized Nash Equilibria (GNE) are often computed using the normalized solution concept, which assumes identical Lagrange multipliers for shared constraints across all players. While widely used, this approach excludes other potentially valuable GNE. This paper presents a novel method based on the Mixed Complementarity Problem (MCP) formulation to compute non-normalized GNE, expanding the solution space. We also propose a systematic approach for selecting the optimal GNE based on predefined criteria, enhancing practical flexibility. Numerical examples illustrate the method’s effectiveness, offering an alternative to traditional normalized solutions.
1 Introduction
@@
Computing a Nash Equilibrium (NE) in a Dynamic Game (DG) is a well-established concept. In DGs, players often share constraints, making each player’s feasible strategy set dependent on others’ strategies. This extends the problem to a Generalized Nash Equilibrium Problem (GNEP). NE serves as a fundamental tool for analyzing competitive behavior across various domains, including economics [1] and racing [2, 3].
@@
In general, the existence or uniqueness of a GNE is not guaranteed. We are interested in problems where GNEPs admit multiple solutions. The standard approach for solving GNEPs is to derive and combine the Karush-Kuhn-Tucker (KKT) conditions for each player’s optimization problem [4]. When shared constraints exist, the Lagrange multipliers associated with these constraints are typically assumed to be identical for all players. This concept, known as the “normalized solution”, was introduced by Rosen [5], who demonstrated its existence and uniqueness under specific conditions. This solution also arises from reformulating the DG as a Quasi Variational Inequality (QVI) problem. Harker [6] and others [7] further showed that this reformulation reduces the solution set to the normalized solution, provided it exists. The normalized solution is widely adopted in various applications, including racing [2, 8, 3], economics [9], and traffic systems [10], and has become the standard approach for solving DGs due to its computational, numerical, and conceptual advantages. Despite these benefits, the normalized solution may not always capture the complexity of real-world interactions, where other GNEs could better represent asymmetric or more intuitive behavior. However, computing non-normalized GNEs is challenging. Standard approaches to solving GNEPs, such as reformulating the KKT conditions as a QVI is inherently bias toward normalized solutions [7]. Reformulating the problem as Mixed Complementarity Problem (MCP) is practically also biased toward the normalized solution by many numerical solvers even when no such constraint is imposed on the solution space. This paper tackles these limitations through two key contributions: 1) Proposing a method that extends the MCP framework to calculate non-normalized GNEs for GNEPs, thereby broadening the solution space, and 2) Introducing a systematic approach for selecting an “optimal” GNE from this expanded solution set, based on predefined criteria.
2 Problem Formulation
@@
Give $N$ players, each player $i\in\{1,...,N\}$ controls the variables $x^{i}\in\mathbb{R}^{n_{i}}$. The vector $x\in\mathbb{R}^{n}$ is formed by concatenating all the players’ decision variables:
$x:=[(x^{1})^{T},...,(x^{N})^{T}]^{T}$ (1)
@@
where $n:=\sum^{N}_{i=1}n_{i}$. $x^{-i}$ represent the decision variables of all players except player $i$. To emphasize the decision variables of player $i$ within $x$ we write $(x^{i},x^{-i})$ instead of $x$. The aim of player $i$ is to choose the $x^{i}$ which minimizes its own cost function $J_{i}(x^{i},x^{-i})$ refer to such decision as “strategy”. The feasible set of the $i$-th decision depends on the strategies of the other players:
@@
$\mathcal{X}_{i}(x_{-i}):=\{x_{i}\in\mathbb{R}^{n_{i}}|h_{i}(x_{i})=0,g_{i}(x_{i})\leq 0,s(x_{i},x_{-i})\leq 0\}$ (3)
@@
$\bar{x}^{i}\in\mathcal{S}_{i}(\bar{x}^{-i})\quad\forall i\in\{1,...,N\}$ (4)
A solution to the GNEP is the GNE. A GNE is the set of strategies that characterized by the fact that none of the players can improve unilaterally it’s cost function by changing its strategy in a feasible direction. Furthermore, it is typical to assume the following assumptions:
@@
    Rosen’s setting [5] - set $\mathcal{X}\subset\mathbb{R}^{n}$ is convex and compact.
@@
$\mathcal{L}_{i}:=J_{i}(x_{i},x_{-i})+\mu_{i}^{T}h_{i}(x_{i})+\lambda_{i}^{T}\cdot g_{i}(x_{i})+\sigma_{i}^{T}\cdot s(x_{i},x_{-i})$ (5)
@@
$\begin{split}&\nabla_{x^{i}}\mathcal{L}_{i}=0,\quad\forall i=1,...,N\\ &0\leq\lambda_{i}\perp g_{i}\leq 0,\quad\forall i=1,...,N\\ &0\leq\sigma_{i}\perp s\leq 0,\quad\forall i=1,...,N\\ &\lambda_{i}\geq 0,\quad\forall i=1,...,N\\ &\sigma_{i}\geq 0,\quad\forall i=1,...,N\\ &h_{i}=0,\quad\forall i=1,...,N\\ \end{split}$ (6)
The following theorem for calculating the GNE as a solution to a GNEP.
@@
[4, 11, 12] Given the assumptions above and that some suitable Constraint Qualification (CQ) holds (e.g., Slater CQ, linear independence CQ,…), then
@@
    If the tuple $(\bar{x},\{\bar{\mu}_{i}\}_{i=1}^{N},\{\bar{\lambda}_{i}\}_{i=1}^{N},\{\bar{\sigma}_{i}\}_{i=1}^{N})$ satisfies the KKT conditions (6), then $\bar{x}$ is a solution of the GNEP.
@@
    If $\bar{x}$ is a solution of the GNEP, then there exists a suitable vectors of multipliers ($(\{\bar{\mu}_{i}\}_{i=1}^{N},\{\bar{\lambda}_{i}\}_{i=1}^{N},\{\bar{\sigma}_{i}\}_{i=1}^{N})$ that satisfy the KKT conditions (6) when $x=\bar{x}$.
2.1 GNE of Car Racing Example
To illustrate the formulation presented in Section 2, we consider a single-step, finite-horizon, discrete-time, general-sum, open-loop dynamic game in a racing scenario involving three cars on a two-lane track. Each car is a player in the generalized Nash equilibrium problem (GNEP), where it controls its velocity along a one-dimensional lane. Car 1 occupies the first lane, while Cars 2 and 3 share the second lane. $x_{i}$ and $v_{i}$ represent the final position and velocity of car $i$, respectively. The game is formulated as follows:
$\min_{x_{1},v_{1}}\quad-x_{1}+x_{2}+\frac{1}{2}v_{1}^{2}$ $\min_{x_{2},v_{2}}\quad-x_{2}+x_{1}+\frac{1}{2}v_{2}^{2}$ $\min_{x_{3},v_{3}}\quad-x_{3}+x_{2}+\frac{1}{2}v_{3}^{2}$ (7)
s.t. s.t. s.t.
$x_{1}=x_{1}(0)+\Delta t\cdot v_{1}$ $x_{2}=x_{2}(0)+\Delta t\cdot v_{2}$ $x_{3}=x_{3}(0)+\Delta t\cdot v_{3}$
$x_{2}\leq x_{3}$ $x_{2}\leq x_{3}$ $x_{2}\leq x_{3}$
$x_{1}(0)=0$ $x_{2}(0)=0.5$ $x_{3}(0)=0.75$
where $\Delta t=1[sec]$. The objective of Car 1 is to advance farther than Car 2 while minimizing control effort. Similarly, Car 2 aims to advance farther than Car 1 while minimizing its control effort. Car 3, however, seeks to assist Car 1 in winning the race while keeping its control effort minimal. A collision avoidance constraint between Cars 2 and 3 makes it possible for car 3 to affect car 2 trajectory. Figure 1 illustrates the initial conditions of the race. Throughout this paper, we will use this example to provide insights into the proposed methods.
@@ 2.1 GNE of Car Racing Example
3 The Normalized Solution For Dynamic Games
Rosen introduced the concept of the “Normalized” solution for the Generalized Nash Equilibrium Problem (GNEP) in [5]. Under the previously stated assumptions, the normalized solution corresponds to a scenario where the Lagrange multipliers associated with the shared constraints are identical for all players:
$\sigma^{1}=\dots=\sigma^{N}=\sigma.$ (8)
Rosen showed that a unique normalized solution exists under the specified conditions. Subsequent works, including [6, 7], showed that reformulating the GNEP as a Quasi-Variational Inequality (QVI) problem inherently reduces the solution set to the normalized solution, see [7] section 3. As a result, the normalized solution has become the de facto standard in solving GNEPs, to the extent that many papers do not explicitly mention this assumption. One popular method for solving GNEPs is to reformulate the Karush-Kuhn-Tucker (KKT) conditions as a Mixed Complementarity Problem (MCP). An MCP is defined for a function $F(z):\mathbb{R}^{d}\to\mathbb{R}^{d}$, with lower and upper bounds $l\in\mathbb{R}^{d}$ and $u\in\mathbb{R}^{d}$, respectively. The goal is to find a vector $z^{*}\in\mathbb{R}^{d}$ such that the following conditions hold for each element $j\in\{1,\dots,d\}$:
@@ 3 The Normalized Solution For Dynamic Games
$F(z)$ $=\begin{bmatrix}\nabla_{x^{1}}\mathcal{L}_{1}\\ \vdots\\ \nabla_{x^{N}}\mathcal{L}_{N}\\ h_{1}\\ \vdots\\ h_{N}\\ -g_{1}\\ \vdots\\ -g_{N}\\ -s\\ \vdots\\ -s\end{bmatrix},\quad l=\begin{bmatrix}-\infty\\ \vdots\\ -\infty\\ -\infty\\ \vdots\\ -\infty\\ 0\\ \vdots\\ 0\\ 0\\ \vdots\\ 0\end{bmatrix},\quad u=\begin{bmatrix}\infty\\ \vdots\\ \infty\\ \infty\\ \vdots\\ \infty\\ \infty\\ \vdots\\ \infty\\ \infty\\ \vdots\\ \infty\end{bmatrix},$ (10)
@@
where $\bar{z}=[x^{1},\dots,x^{N},\mu_{1},\dots,\mu_{N},\lambda_{1},\dots,\lambda_{N},\sigma_{1},\dots,\sigma_{N}]$ represents the vector of all decision variables and Lagrange multipliers, and the shared constraint $s$ is duplicated $N$ times. Solving the MCP above yields the normalized solution, as shown in [7]. To simplify the formulation, the normalized solution assumption reduces the MCP by enforcing shared multipliers resulting in:
$F(z)$ $=\begin{bmatrix}\nabla_{x^{1}}\mathcal{L}_{1}\\ \vdots\\ \nabla_{x^{N}}\mathcal{L}_{N}\\ h_{1}\\ \vdots\\ h_{N}\\ -g_{1}\\ \vdots\\ -g_{N}\\ -s\end{bmatrix},\quad l=\begin{bmatrix}-\infty\\ \vdots\\ -\infty\\ -\infty\\ \vdots\\ -\infty\\ 0\\ \vdots\\ 0\\ 0\end{bmatrix},\quad u=\begin{bmatrix}\infty\\ \vdots\\ \infty\\ \infty\\ \vdots\\ \infty\\ \infty\\ \vdots\\ \infty\\ \infty\end{bmatrix},$ (11)
where $z=[x^{1},\dots,x^{N},\mu_{1},\dots,\mu_{N},\lambda_{1},\dots,\lambda_{N},\sigma]$ and $\sigma^{1}=\cdots=\sigma^{N}=\sigma$.
@@
Numerical solvers, such as PATH [13], are widely used to solve MCPs efficiently and are considered state-of-the-art tools for this purpose. The normalized solution remains a computationally efficient and conceptually straightforward approach for solving GNEPs, despite its limitations in representing more complex dynamics.
3.1 Limitations of Normalized GNE solutions

### v3

# Non-Normalized Solutions of Generalized Nash Equilibria in Dynamic Games With Shared Constraints

@@ Abstract
In dynamic games with shared constraints, Generalized Nash Equilibria (GNE) are often computed using the normalized solution concept, which assumes identical Lagrange multipliers for shared constraints across all players. While widely used, this approach excludes other potentially valuable GNE. This paper presents a simple and novel method that can utilize the Mixed Complementarity Problem (MCP) formulation to compute non-normalized GNE. The proposed approach allows to expand the solution space of GNE. We also propose a systematic approach for selecting one GNE based on a predefined optimality criteria, enhancing practical flexibility. Numerical examples from past literature and a car racing example illustrate the method’s effectiveness, offering an alternative to traditional normalized solutions.
I Introduction
@@
Computing a Nash Equilibrium (NE) in a Dynamic Game (DG) is a well-established concept. In DGs, players often share constraints, making each player’s feasible strategy set dependent on others’ strategies. This extends the problem to a Generalized Nash Equilibrium Problem (GNEP). NE serves as a fundamental tool for analyzing competitive behavior across various domains, including economics [arrow2024existence] and competitive sports like racing [zhu2023sequential, spica2020real].
@@
In general, the existence or uniqueness of a GNE is not guaranteed. We are interested in problems where GNEPs admit multiple solutions. The standard approach for solving GNEPs is to derive and combine the Karush-Kuhn-Tucker (KKT) conditions for each player’s optimization problem [dreves2011solution]. When shared constraints exist, the Lagrange multipliers associated with these constraints are typically assumed to be identical for all players. This concept, known as the “normalized solution”, was introduced by Rosen [rosen1965existence], who demonstrated its existence and uniqueness under specific conditions. This solution also arises from reformulating the DG as a Quasi Variational Inequality (QVI) problem. Harker [harker1991generalized] and others [facchinei2007generalized] further showed that this reformulation reduces the solution set to the normalized solution, provided it exists. The normalized solution is widely adopted in various applications, including racing [zhu2023sequential, liu2023learning, spica2020real], economics [bacsar1998dynamic], and traffic systems [migot2019revisiting], and has become the standard approach for solving DGs due to its computational, numerical, and conceptual advantages. Despite these benefits, the normalized solution may not always capture the complexity of real-world interactions, where other GNEs could better represent asymmetric or more intuitive behavior. However, computing non-normalized GNEs is challenging. Standard approaches to solving GNEPs, such as reformulating the KKT conditions as a QVI is inherently bias toward normalized solutions [facchinei2007generalized]. Reformulating the problem as Mixed Complementarity Problem (MCP) is practically also biased toward the normalized solution by many numerical solvers even when no such constraint is imposed on the solution space. The solution by the iterative best response [spica2020real, wang2021game] may result in a non-normalized solution but any solution achieved by this method is heavily dependent on the initial guess and convergence is not guaranteed. The work done by Nabetani et al [nabetani2011parametrized] suggest a method to calculate Non-Normalized solution by reformulating the QVI in a way that de-normalizes the solution. The resulting method requires solving a large number of VI instances - Each instance may or may not correspond to a valid GNE, even if the VI solver converges. For linear generalized Nash equilibrium problems, Dreves [dreves2017computing] developed a duality-based algorithm that computes the entire solution set, not only normalized equilibria, and proved finite termination.
This paper addresses the limitations of existing GNEP solution methods through three key contributions:
  1. 1.
    A novel MCP-based framework for non-normalized equilibria: We extend the Mixed Complementarity Problem (MCP) formulation to systematically compute non-normalized Generalized Nash Equilibria (GNE), thereby uncovering solutions that are typically excluded by standard formulations.
  2. 2.
    A GNE selection mechanism: We propose a systematic bi-level optimization approach for selecting an “optimal” equilibrium from the expanded set of non-normalized GNEs. Crucially, we formally show that the introduced scaling factors act as well-defined tuning knobs for navigating and parameterizing the entire GNE solution set.
  3. 3.
    Empirical validation: We demonstrate both analytically and numerically how normalized solutions can lead to unrealistic or counterintuitive outcomes, and we show that the proposed framework not only overcomes these issues but also yields richer, more intuitive, and practically relevant equilibria.
The paper is organized as follows. Section II introduces the problem formulation and a motivating example. Section III reviews the classical normalized GNE approach and its limitations. Section IV presents the proposed MCP-based method for computing non-normalized GNEs and the theoretical results that establish the scaling matrices as formal tuning parameters. Section V provides analytical validation, while Section VI introduces the bi-level optimization for equilibrium selection. Section VII presents simulation results on dynamic racing games, and Section VIII concludes the paper.
II Problem Formulation
@@
Given $M$ players, each player $i\in\{1,...,M\}$ controls the variables $x^{i}\in\mathbb{R}^{n_{i}}$. The vector $x\in\mathbb{R}^{n}$ is formed by concatenating all the players’ decision variables:
$x:=[(x^{1})^{T},...,(x^{M})^{T}]^{T}$ (1)
@@
where $n:=\sum^{M}_{i=1}n_{i}$. $x^{-i}$ represent the decision variables of all players except player $i$. To emphasize the decision variables of player $i$ within $x$ we write $(x^{i},x^{-i})$ instead of $x$. The aim of player $i$ is to choose the $x^{i}$ which minimizes its own cost function $J_{i}(x^{i},x^{-i})$ refer to such decision as “strategy”. The feasible set of the $i$-th decision depends on the strategies of the other players:
@@
$\begin{split}\mathcal{X}_{i}(x_{-i}):=\{x_{i}\in\mathbb{R}^{n_{i}}|h_{i}(x_{i})&=0\\ ,g_{i}(x_{i})&\leq 0,~s(x_{i},x_{-i})\leq 0\}\end{split}$ (3)
@@
$\bar{x}^{i}\in\mathcal{S}_{i}(\bar{x}^{-i})\quad\forall i\in\{1,...,M\}$ (4)
A solution to the GNEP is the GNE. A GNE is the set of strategies that characterized by the fact that none of the players can improve unilaterally its cost function by changing its strategy in a feasible direction. Furthermore, it is typical to assume the following assumptions (Rosen’s setting [rosen1965existence]):
Assumption 1.
@@
    set $\mathcal{X}\subset\mathbb{R}^{n}$ is convex and compact.
@@
$\mathcal{L}_{i}:=J_{i}(x^{i},x^{-i})+\mu_{i}^{T}h_{i}(x^{i})+\lambda_{i}^{T}\cdot g_{i}(x^{i})+\sigma_{i}^{T}\cdot s(x^{i},x^{-i})$ (5)
@@
[…]
The following theorem summarizes known results on computing GNE as a solution to a GNEP.
@@
[dreves2011solution, bueno2019optimality, facchinei2010generalized] Consider the game (2)-(3), assume Assumption 1 and some suitable Constraint Qualification (e.g., Slater Conditions) hold, then
@@
    If the tuple $(\bar{x},\{\bar{\mu}_{i}\}_{i=1}^{M},\{\bar{\lambda}_{i}\}_{i=1}^{M},\{\bar{\sigma}_{i}\}_{i=1}^{M})$ satisfies the KKT conditions (6), then $\bar{x}$ is a solution of the GNEP.
@@
    If $\bar{x}$ is a solution of the GNEP, then there exists a suitable vectors of multipliers $(\{\bar{\mu}_{i}\}_{i=1}^{M},\{\bar{\lambda}_{i}\}_{i=1}^{M},\{\bar{\sigma}_{i}\}_{i=1}^{M})$ that satisfy the KKT conditions (6) when $x=\bar{x}$
  3. 3.
    If $\bar{x}$ is a solution of the GNEP and LICQ holds, then there exists a unique vectors of multipliers $(\{\bar{\mu}_{i}\}_{i=1}^{M},\{\bar{\lambda}_{i}\}_{i=1}^{M},\{\bar{\sigma}_{i}\}_{i=1}^{M})$ that satisfy the KKT conditions (6) when $x=\bar{x}$
II-A GNE Modeling of a Three Car Racing Problem
To illustrate the formulation presented in Section II, we consider a one-step, finite-horizon, discrete-time, general-sum, open-loop dynamic game in a racing scenario involving three cars on a two-lane track. Each car is a player in the generalized Nash equilibrium problem (GNEP), which controls its velocity along a one-dimensional lane. Car 1 occupies the first lane, while Cars 2 and 3 share the second lane (Figure 1). $x_{i}$ and $v_{i}$ represent the final position and velocity of the car $i$, respectively. The game for player $i\in\{1,2,3\}$:
$\min_{x_{i},v_{i}}J_{i}(x_{i},v_{i},x^{-i})$ (7)
s.t.
$x_{i}=x_{i}(0)+v_{i}\cdot\Delta t$
$x_{2}\leq x_{3}$
[…]
where $\Delta t=1[sec]$. The cost functions of the players are:
$J_{1}=-x_{1}+x_{2}+\frac{1}{2}v_{1}^{2}$ (8)
$J_{2}=-x_{2}+x_{1}+\frac{1}{2}v_{2}^{2}$
$J_{3}=-x_{1}+x_{2}+\frac{1}{2}v_{3}^{2}$
The objective of Car 1 is to advance farther than Car 2 while minimizing control effort. Similarly, Car 2 aims to advance farther than Car 1 while minimizing its control effort. Car 3, however, seeks to assist Car 1 in winning the race while keeping its control effort minimal. A collision avoidance constraint between Cars 2 and 3 makes it possible for car 3 to affect car 2 position. Figure 1 illustrates the initial conditions of the race. Throughout this paper, we will use this example to provide insights into the proposed methods.
@@ II-A GNE Modeling of a Three Car Racing Problem
III The Normalized Solution For Dynamic Games
Rosen introduced the concept of the “Normalized” solution for the Generalized Nash Equilibrium Problem (GNEP) in [rosen1965existence]. Under the previously stated assumptions, the normalized solution corresponds to a scenario where the Lagrange multipliers associated with the shared constraints are identical for all players:
$\sigma^{1}=\dots=\sigma^{M}=\sigma.$ (9)
Rosen showed that a unique normalized solution exists under the specified conditions. Subsequent works, including [harker1991generalized, facchinei2007generalized], showed that reformulating the GNEP as a Quasi-Variational Inequality (QVI) problem inherently reduces the solution set to the normalized solution, see [facchinei2007generalized] section 3. As a result, the normalized solution has become the de facto standard in solving GNEPs, to the extent that many papers do not explicitly mention this assumption. One popular method for solving GNEPs is to reformulate the Karush-Kuhn-Tucker (KKT) conditions as a Mixed Complementarity Problem (MCP). An MCP is defined for a function $F(z):\mathbb{R}^{d}\to\mathbb{R}^{d}$, with lower and upper bounds $l\in\mathbb{R}^{d}$ and $u\in\mathbb{R}^{d}$, respectively. The goal is to find a vector $z^{*}\in\mathbb{R}^{d}$ such that the following conditions hold for each element $j\in\{1,\dots,d\}$:
@@ III The Normalized Solution For Dynamic Games
$\begin{split}F(z)&=[\nabla^{T}_{x^{1}}\mathcal{L}_{1},\cdots,\nabla^{T}_{x^{M}}\mathcal{L}_{M},h^{T}_{1},\cdots,h^{T}_{M},\cdots\\ &\qquad\qquad\qquad\cdots,-g^{T}_{1},\cdots,-g^{T}_{M},-s^{T},\cdots,-s^{T}]^{T}\\ l&=[-\infty,\cdots,-\infty,-\infty,\cdots,-\infty,0,\cdots,0,0,\cdots,0]^{T}\\ u&=[\infty,\cdots,\infty,\infty,\cdots,\infty,\infty,\cdots,\infty,\infty,\cdots,\infty]^{T}\end{split}$ (11)
@@
where $\bar{z}=[x^{1},\dots,x^{M},\mu_{1},\dots,\mu_{M},\lambda_{1},\dots,\lambda_{M},\sigma_{1},\dots,\sigma_{M}]$ represents the vector of all decision variables and Lagrange multipliers, and the shared constraint $s$ is duplicated $N$ times. Solving the MCP above yields the normalized solution, as shown in [facchinei2007generalized]. To simplify the formulation, the normalized solution assumption reduces the MCP by enforcing shared multipliers resulting in:
$\begin{split}F(z)&=[\nabla^{T}_{x^{1}}\mathcal{L}_{1},\cdots,\nabla^{T}_{x^{M}}\mathcal{L}_{M},h^{T}_{1},\cdots,h^{T}_{M},\cdots\\ &\qquad\qquad\qquad\qquad\cdots,-g^{T}_{1},\cdots,-g^{T}_{M},-s^{T}]^{T}\\ l&=[-\infty,\cdots,-\infty,-\infty,\cdots,-\infty,0,\cdots,0,0]^{T}\\ u&=[\infty,\cdots,\infty,\infty,\cdots,\infty,\infty,\cdots,\infty,\infty]^{T}\end{split}$ (12)
where $z=[x^{1},\dots,x^{M},\mu_{1},\dots,\mu_{M},\lambda_{1},\dots,\lambda_{M},\sigma]$.
@@
Numerical solvers, such as PATH [dirkse1995path], are widely used to solve MCPs efficiently and are considered state-of-the-art tools for this purpose. The normalized solution remains a computationally efficient and conceptually straightforward approach for solving GNEPs, despite its limitations in representing more complex dynamics.
III-A Limitations of Normalized GNE solutions